OpenStudy (anonymous):

Put in point slope form and slope intercept form y1=14% y2=12.5% x1=5 x2=13 I don't know how to find the slope

zepdrix (zepdrix):

Hey Britt c: What are the percents about? 0_o Hmm

OpenStudy (anonymous):

A multi-year analysis was begun in 1987 to determine the percentage of American women claiming to be on a weight-loss diet. 5 years into the study, 14\% of women were found to be on a diet, while 13 years into the study, 12.5\% of women were found to be on a diet

zepdrix (zepdrix):

Hmm ok good, so you've labeled your x's and y's correctly :)

OpenStudy (anonymous):

Let \, y \, = percentage of women on a diet, and let \, x \, = number of years into the study

OpenStudy (anonymous):

yes :)

zepdrix (zepdrix):

Taking the different of the y's tell us the `vertical height` between the points. \(\Large\rm y_2-y_1\) we often call this "rise". Taking the difference of the x's tells us the `horizontal distance` between the points. \(\Large\rm x_2-x_1\) we often call this "run". Our slope is the ratio of the rise to the run. Change in height divided by change in ... distance or whatever. \[\Large\rm slope=\frac{y_2-y_1}{x_2-x_1}\] So we just need to plug in those points and simplify :O

zepdrix (zepdrix):

You're using Point-Slope form of a line? Ooo goodie, that mean you'll have a little bit less work. We need only find the slope, then plug the stuff in.

zepdrix (zepdrix):

\[\Large\rm slope=\frac{12.5\text{%}-14\text{%}}{13-5}\]Understand how I plugged those into the formula? :O

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Simplify the top and bottom! c: What do you get?

zepdrix (zepdrix):

Oh boy :( The website is tweaking out again!

OpenStudy (anonymous):

i got -.015/8

OpenStudy (anonymous):

but i can't use decimals

zepdrix (zepdrix):

You can't? :O Oh I see.

zepdrix (zepdrix):

What happened in your numerator? Ohhh did you turn the % into a decimal or something?

zepdrix (zepdrix):

\[\Large\rm slope=\frac{-1.5\text%}{8}\]and you wrote it as,\[\Large\rm slope=\frac{-.015}{8}\]Umm I guess to get rid of the decimal, we'll have to multiply by a bunch of 10's. Multiplying by 10 moves the decimal one place. Multiplying by two tens (which is the same as multiplying by 100) will move it two places. So we'll need to multiply our expression, top and bottom, by 1000, to move the decimal three places.\[\Large\rm slope=\frac{-.015}{8}\left(\frac{1000}{1000}\right)\]

zepdrix (zepdrix):

\[\Large\rm slope=-\frac{150}{8000}\]

zepdrix (zepdrix):

This number doesn't make a whole lot of sense though D: It's telling us that 150 less people will be on this diet every 8000 years. -_- Can't we just leave it in percent form or no? :o

OpenStudy (anonymous):

umm let me see

OpenStudy (anonymous):

i'm not sure. this is the problem attached

zepdrix (zepdrix):

MSU? :D Oooo neato

zepdrix (zepdrix):

Did you misread the question maybe silly? :D I see a line that says "Decimals `are` allowed throughout this problem".

OpenStudy (anonymous):

oh whoops, i just assumed because all the other ones it said no decimals okay hold on ha

zepdrix (zepdrix):

ah :)

OpenStudy (anonymous):

is it -.015/8(x-.125)?

zepdrix (zepdrix):

Woops your `x` isn't plugged in correctly. You plugged your y2 into it.

zepdrix (zepdrix):

I would simplify your slope a little bit further. If you want to leave it as an ugly fraction like that it's fine.... But if you just toss it into the calculator it gives you: slope=-0.1875 ya?

OpenStudy (anonymous):

-0.1875(x-5) plugged it in and its correct :)

zepdrix (zepdrix):

yay team \c:/

OpenStudy (anonymous):

hmm but i tried to put it in slope intercept form and got .1875x+14.94 and it says its wrong

zepdrix (zepdrix):

Hmm slope intercept form is a little tricky D: We can get there from point-slope form though.

zepdrix (zepdrix):

\[\Large\rm y-14=-.1875(x-5)\]Distributing the -.1875 gives us,\[\Large\rm y-14=-.1875x+0.9375\]Adding 14,\[\Large\rm y=-.1875x+14.9375\]WHUT? It looks like you did it correctly. Wrong answer? D: Is the rounding the problem maybe?